OpenStudy (anonymous):

***HELP*** i think its slope form Need help with steps to solve: y=10x+7

OpenStudy (amistre64):

your question is vague. what is it you need help doing?

OpenStudy (anonymous):

i need help with the steps to do this equation. i did it a while back. now i cant seem to be able to figure out how to do it.

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

what is it that you are spose to be 'doing'? it seems like you are trying to get the line equation into some format ... but what format?

OpenStudy (anonymous):

the formula is y=mx+b

OpenStudy (anonymous):

i know the answer to it, i just need help explaining the steps.

OpenStudy (amistre64):

thats not a formula, thats a format for a line in slope intercept form.

OpenStudy (amistre64):

are you trying to find the slope and the intercept?

OpenStudy (anonymous):

yes, i need the steps to fiding both. please!

OpenStudy (amistre64):

given: y = mx + b this is called the 'slope intercept' form simply because it uses the slope (m) and the y intercept (b) in the construction of the equation:

OpenStudy (anonymous):

okay. but that doesn't teach me the steps to do it..

OpenStudy (amistre64):

the steps are just looking at it. then picking the m and the b values

OpenStudy (anonymous):

Danielle is playing the newest app on her phone, Algebra City Defense. Earlier in the game, you and Danielle helped the Mathonauts defend the city against a giant, three-headed monster. After defeating the monster, the game advances to Level Two! You and Danielle must help the Mathonauts blast off to their Interstellar Headquarters. Before blast-off, the Mathonauts must set the trajectory of their ship. Create a linear equation with a positive slope to be your original trajectory. After blast-off, the sensors have picked up an incoming meteor. The Mathonauts frantically start preparing for impact. If the linear equation of the meteor’s path is known, describe to the captain how to solve where your equation from question 1 and the meteor’s path will cross. Explain any possible methods used in discovering a solution. Having survived the meteor impact, thanks to some last-minute evasive maneuvers, the Mathonauts now set their sights on their Interstellar Headquarters. The Interstellar Headquarters orbits the Earth based on the equation y2 + x2 = 40,000. Using the original trajectory of the ship and complete sentences, explain to the pilot how to find where the ship’s path will cross the Interstellar Headquarters’s path. While docked, the captain wants to resupply the ship’s missiles and ray-guns. You will need to explain to the captain what information you must know and how that can be used to determine the number of missiles and ray-guns to fill the ship and use all the batteries. The ship can hold only 15 weapons total and only has 45 batteries. The missiles and ray-guns each require different amounts of batteries to operate. Before departing the Interstellar Headquarters, the ship’s navigator begins to plan the next trip. She uses a map with their current trajectory already graphed on it. There is also a table that has coordinate points of a satellite that the Mathonauts must intercept. Explain to the navigator how she can use the graph and table to find where they will intercept the satellite. Assume the path of the satellite is linear. Use the coordinates to create an equation representing the path of the satellite. Explain the process and show how to find the point of intersection algebraically. This is what im doing...

OpenStudy (amistre64):

if you have a different format, and are asked to determine the slope and y intercept, then one approach is to 'solve for y' so that you convert it into a 'slope intercept' format. but as is, you have the slope (m) defined already in the equation, and the y intercept (b). all it takes is to look at the values that they use and write them down :/

OpenStudy (amistre64):

sounds like a bunch of system of equation stuff to toy with

OpenStudy (anonymous):

@amistre64 can u help me by explaiing it better.

OpenStudy (amistre64):

Before blast-off, the Mathonauts must set the trajectory of their ship. Create a linear equation with a positive slope to be your original trajectory. assuming your equation is starting at 0,0 what trajectory did you pick?

OpenStudy (amistre64):

and now the site is giving me connection errors now so this doesnt look to promising ...

OpenStudy (anonymous):

y=10x+7

OpenStudy (anonymous):

meaning u dunn thin u will be able to help me?

OpenStudy (amistre64):

it means that the site is starting to misbehave for me and will most likely freeze up soon is all

OpenStudy (amistre64):

the next part is simply to tell what youve learned so far about finding where 2 lines cross. what have you learned so far?

OpenStudy (anonymous):

oh, okay. darn.

OpenStudy (anonymous):

um, i honestly cant remember anything.. D:

OpenStudy (amistre64):

graphing is the most basic approach and usually the most unreliable since its based on more of an estimation, unless a program maps it out and gives you the intersection point. other methods involve substitutions, eliminations, and perhaps matrix solutions.

OpenStudy (anonymous):

i am terrible with graphing.. i think the subtraction one is the one i used last time.

OpenStudy (amistre64):

the circle one is ideal for substitution since we already have a y value to play with: x^2 + y^2 = r^2, but we know y = 10x+7 x^2 + (10x+7)^2 = r^2 this then turns it into a quadratic equation that can be solved by othe rmethods

OpenStudy (anonymous):

let me see if i can solve it.

OpenStudy (anonymous):

nope.. all the letters in the equation are confusing me..

OpenStudy (amistre64):

substitution is useful since it tends to convert a 2 variable setup, into a 1 variable setup. solving for 1 variable is simpler than finding 2

OpenStudy (amistre64):

r^2 is defined for us, i just used r^2 since its simpler to type

OpenStudy (anonymous):